# Steinmetz Neural Networks for Complex-Valued Data

### v2

@@ 5 Analytic Neural Network
As detailed in Section 4, by introducing a constraint within the latent representation, $Z^{m}$, such that $Z_{R}^{m}$ and $Z_{I}^{m}$, are related by a deterministic, bijective function, $\phi(\cdot)$, we can leverage improved control over the generalization error, $\Delta s$. A natural question that follows is: which $\phi(\cdot)$ should be chosen to improve predictive performance? To address this, we consider a configuration that focuses on the properties and predictive advantages of orthogonal latent representations.
@@
Consider $\{x_{R}^{m},m\in\mathcal{V}\}$ and $\{x_{I}^{m},m\in\mathcal{V}\}$ from the training dataset, with $x_{R}^{m},x_{I}^{m}\in\mathbb{R}^{dN}$. It follows that $z_{R}^{m}=g(x_{R}^{m})$ and $z_{I}^{m}=f(x_{I}^{m})$, wherein $\smash{z_{R}^{m},z_{I}^{m}\in\mathbb{R}^{lN}}$. To implement the Hilbert consistency penalty, we make use of the discrete Fourier transform (DFT), leveraging its properties in relation to phase shifts — we consider $F_{R}^{m}=\mathcal{F}\{z_{R}^{m}\}\in\mathbb{C}^{lN}$ as the DFT of $z_{R}^{m}$ where $b$ denotes the frequency index. Eq. (32) summarizes the frequency domain implementation of this phase shift.
@@
Table 1: Test performance on CV-MNIST $(M=500)$, CV-CIFAR-10 $(M=50{,}000)$, RASPNet $(M=20{,}000)$.
@@
(c) RASPNet
Figure 3: Test performance on CV-MNIST $(M=500)$, CV-CIFAR-10 $(M=50{,}000)$, and RASPNet $(M=20{,}000)$ using CVNN, RVNN, Steinmetz neural network, and analytic neural network. The x-axis represents the training epochs, while the y-axis indicates the test performance (classification accuracy and mean squared error).
@@ 6 Empirical Results
We present empirical results on benchmark datasets for complex-valued multi-class classification and regression, and on a synthetic signal processing example for complex-valued regression. Per (Trabelsi et al., 2018) and (Scardapane et al., 2020), we present the classification results on complex-valued MNIST (Deng, 2012) and CIFAR10 (Krizhevsky et al., 2009). We also present a regression result on the RASPNet benchmark dataset (Venkatasubramanian et al., 2024), which comprises intrinsically complex-valued radar returns. The CVNNs in this analysis were constructed using the Complex Pytorch library (Matthès et al., 2021), which implements the layers proposed in (Trabelsi et al., 2018).
6.1 Benchmark Datasets
This first experiment provides an assessment of our methods on Complex-Valued MNIST (CV-MNIST) and Complex-Valued CIFAR-10 (CV-CIFAR-10) for multi-class classification, and on RASPNet for localization, evaluating the efficacy of the proposed Steinmetz and analytic neural networks when there are no ablations introduced within the data. We obtain CV-MNIST by taking a $dN=784$-point DFT of each of the $M=500$ training images (a small subset comprising the first $500$ images in MNIST). The test set is formed by taking a $784$-point DFT of each of the $10{,}000$ test images. We similarly obtain CV-CIFAR-10 by taking a $dN=3072$-point DFT of each of the $M=50{,}000$ training images. The test set is formed by taking a $3072$-point DFT of each of the $10{,}000$ test images. For both CV-MNIST and CV-CIFAR-10, we have $k=10$. RASPNet consists of $M=20{,}000$ training and $5{,}000$ test radar returns of size $dN=1680$, with $k=3$ ($3$D target localization).
@@ 6.1 Benchmark Datasets
Figure 4: Noise robustness test performance on CV-MNIST ($M=60{,}000$) and CV-CIFAR-10 ($M=50{,}000$) using CVNN, RVNN, Steinmetz neural network, and analytic neural network. The x-axis is the scaling factor, $\eta$, for the additive complex normal noise, while the y-axis indicates the classification accuracy.
@@
We train a RVNN, CVNN, Steinmetz neural network, and analytic neural network using Cross Entropy Loss for $lN=64$ to classify the images in CV-MNIST and CV-CIFAR-10, and using MSE Loss for $lN=128$ to localize targets from the radar returns in RASPNet. These neural network architectures and hyperparameter choices are described in Section D of the Appendix, wherein we use the Adam optimizer (Kingma and Ba, 2014) to train each architecture using a fixed learning rate. The empirical results pertaining to this first experiment are depicted in Table 1 and Figure 3. On CV-MNIST, CV-CIFAR-10, and RASPNet, we observe that the Steinmetz and analytic neural networks achieve improved generalization over the classical RVNN and CVNN. The analytic neural network achieves the highest classification accuracy and lowest MSE.
@@ 6.2 Channel Identification
We evaluate our proposed Steinmetz and analytic networks on the benchmark channel identification task from (Scardapane et al., 2020; Bouboulis et al., 2015). Let $\smash{X^{m}=\sqrt{1-\rho^{2}}\bar{X}\vphantom{X}^{m}+i\rho\tilde{X}\vphantom{X}^{m}}$ denote the input to the channel, wherein $\smash{\bar{X}\vphantom{X}^{m}}$ and $\smash{\tilde{X}\vphantom{X}^{m}}$ are Gaussian random variables, and $\rho$ determines the circularity of the signal. Here, $dN=5$ denotes the length of the input sequence, an embedding of the channel inputs over $dN$ time steps. The channel output, $Y^{m}\in\mathbb{C}^{k}$, is formed using a linear filter, a memoryless nonlinearity, and by adding white Gaussian noise to achieve an SNR of $5$ dB, where $X^{m}$ and $Y^{m}$ are equivalent to $s_{n}$ and $r_{n}$ from (Scardapane et al., 2020). We consider $M=1000$ training examples and $1000$ test examples. Each of the real-valued architectures output a $22k=$ 2D vector, $\smash{[[\hat{Y}_{R}^{m}]^{T},[\hat{Y}_{I}^{m}]^{T}]^{T}}$, where $\smash{\hat{Y}_{R}^{m},\hat{Y}_{I}^{m}\in\mathbb{R}}$, and the CVNN outputs a $k=1$-dimensional scalar, $\smash{\hat{Y}^{m}=\hat{Y}_{R}^{m}+i\hat{Y}_{I}^{m}}$, which we reshape to form $\smash{[[\hat{Y}_{R}^{m}]^{T},[\hat{Y}_{I}^{m}]^{T}]^{T}}$. We train our Steinmetz and analytic neural networks, RVNN, and CVNN using MSE Loss, minimizing the distance between $\smash{[[\hat{Y}_{R}^{m}]^{T},[\hat{Y}_{I}^{m}]^{T}]^{T}}$ and $\smash{[[Y_{R}^{m}]^{T},[Y_{I}^{m}]^{T}]^{T}}$. We select $lN=64$, and use the Adam optimizer to train each architecture using a fixed learning rate. We compute and report the MSE between the predicted and true magnitudes and phases on the test dataset in Table 2. From Table 2, we see while the magnitude prediction error is comparable between the CVNN, RVNN, and the Steinmetz and analytic neural networks, the latter pair observes a much lower phase prediction error.
@@ 7 Conclusion
In this work, we introduced Steinmetz neural networks, a new approach to processing complex-valued data using DNNs with parallel real-valued subnetworks. We provided its mathematical framework and outlined a consistency constraint to lower its generalization error upper bound, and we presented the analytic neural network, which incorporates the consistency penalty, for practical implementation. We evaluated these networks on regression and classification benchmarks, showing improvements over existing RVNNs and CVNNs. Future work includes investigating more effective training techniques for Steinmetz neural networks and theoretical performance guarantees.
@@ Appendix A Consistency Constraint Derivation
As outlined in Section 4, our aim is to exploit the Steinmetz neural network architecture by deriving a consistency constraint that allows for improved control over the generalization error, $\Delta s$. Recall Eq. 4.1, which provides an upper bound on $\Delta s$ in terms of the mutual information between $X^{m}$ and $Z^{m}$, between $Y^{m}$ and $Z^{m}$, and between $S$ and $\theta_{\psi_{S}}$. The complete derivation of this bound is provided in [Kawaguchi et al., 2023].
@@
For improved control over the generalization error, we form a smaller upper bound on $\Delta s$ by deriving a lower bound, $\mathcal{D}(Z^{m})$, on $[I(X^{m};Z^{m})-I(Y^{m};Z^{m})+I(S;\theta_{\psi_{S}})]$. We previously stated $\mathcal{D}(Z^{m})$ is achievable by imposing a constraint on the latent representation, $Z^{m}$ (see Corollary 4.2).
@@
This result is also summarized in Theorem 4.3 of the main text. We extend this result to derive the smaller upper bound on the generalization error, $\Delta s$, provided in Theorem 4.4 of the main text.
@@ C.2 CV-CIFAR-10 Dataset
C.3 RASPNet Dataset
The RASPNet dataset for radar adaptive signal processing consists of radar returns gathered from $100$ realistic radar scenarios located in the contiguous United States. We consider the scenario index $i=29$ from RASPNet, which corresponds to the Bonneville Salt Flats, UT, and consider the real and imaginary parts of a size $21\times 16\times 5$ radar return, comprising $21$ range bins, $16$ channels, and $5$ realizations (single pulse transmission). Accordingly, each sample is defined by a $dN=1680$-dimensional real feature vector, and a $1680$-dimensional imaginary feature vector. Each radar return contains a single point target, with position encoded in normalized (range, azimuth, elevation) indices. For the regression result from Section 6.1, we consider $M=20{,}000$ training samples and $5{,}000$ test samples from scenario $i=29$, which can be accessed here. The features and labels are summarized as follows:
@@ C.3 RASPNet Dataset
  • Target Variable: The normalized (range, azimuth, elevation) target location ($k=3$).

### v3

@@ 5 Analytic Neural Network
As detailed in Section 4, by introducing a constraint within the latent representation, $Z^{m}$, such that $Z_{R}^{m}$ and $Z_{I}^{m}$, are related by a deterministic, bijective function, $\phi(\cdot)$, we can leverage improved control over the generalization gap, $\Delta s$. A natural question that follows is: which $\phi(\cdot)$ should be chosen to improve predictive performance? To address this, we consider a configuration that focuses on the properties and predictive advantages of orthogonal latent representations.
@@
Consider sets $\{x_{R}^{m},\,m\in\mathcal{V}\}$ and $\{x_{I}^{m},\,m\in\mathcal{V}\}$ from the training dataset, with $x_{R}^{m},x_{I}^{m}\in\mathbb{R}^{dN}$. It follows that $z_{R}^{m}=g(x_{R}^{m})$ and $z_{I}^{m}=f(x_{I}^{m})$, wherein $\smash{z_{R}^{m},z_{I}^{m}\in\mathbb{R}^{lN}}$. To implement the Hilbert consistency penalty, we make use of the discrete Fourier transform (DFT), leveraging its properties in relation to phase shifts — we consider $F_{R}^{m}=\mathcal{F}\{z_{R}^{m}\}\in\mathbb{C}^{lN}$ as the DFT of $z_{R}^{m}$ where $b$ denotes the frequency index. Eq. (32) summarizes the frequency domain implementation of this phase shift.
@@
Table 1: Test performance comparison on CV-MNIST $(M=500)$, CV-CIFAR-10 $(M=50{,}000)$, CV-CIFAR-100 $(M=50{,}000)$, CV-FSDD $(M=2{,}700)$, and RASPNet $(M=20{,}000)$.
@@
(c) Complex-Valued CIFAR-100
(d) Complex-Valued FSDD
(e) RASPNet Scenario $i=29$
Figure 3: Test performance comparison on CV-MNIST $(M=500)$, CV-CIFAR-10 $(M=50{,}000)$, CV-CIFAR-100 $(M=50{,}000)$, CV-FSDD $(M=2{,}700)$, and RASPNet $(M=20{,}000)$ through CVNN, RVNN, Steinmetz neural network, and analytic neural network. The x-axis indicates the training epochs, while the y-axis indicates the test performance (classification accuracy and mean squared error).
@@ 6 Empirical Results
We present empirical results on benchmark datasets for complex-valued multi-class classification and regression, and on a synthetic signal processing example for complex-valued regression. Per (Trabelsi et al., 2018) and (Scardapane et al., 2020), we present classification results on complex-valued MNIST (Deng, 2012), CIFAR-10 (Krizhevsky et al., 2009), CIFAR-100 (Krizhevsky et al., 2009), and FSDD (Jackson et al., 2018). We also present a regression result on the RASPNet benchmark dataset (Venkatasubramanian et al., 2024), which comprises intrinsically complex-valued radar returns. The CVNNs in this analysis were constructed through the Complex Pytorch library (Matthès et al., 2021), which implements the layers proposed in (Trabelsi et al., 2018).
@@
Figure 4: Noise robustness test performance on CV-MNIST ($M=60{,}000$) and CV-CIFAR-10 ($M=50{,}000$) through CVNN, RVNN, Steinmetz neural network, and analytic neural network. The x-axis is the scaling factor, $\eta$, for the additive complex normal noise, while the y-axis indicates the classification accuracy.
@@
6.1 Benchmark Datasets
The first experiment we consider is an assessment of our methods on Complex-Valued MNIST (CV-MNIST), Complex-Valued CIFAR-10 (CV-CIFAR-10), Complex-Valued CIFAR-100 (CV-CIFAR-100), and the audio-based Complex-Valued Free Spoken Digit Dataset (CV-FSDD), for multi-class classification, and on RASPNet for regression, evaluating the efficacy of the proposed Steinmetz and analytic neural networks when there are no ablations introduced within the data. We obtain CV-MNIST by taking a $dN=784$-point DFT of each of the $M=500$ training images (a small subset of the first $500$ images in MNIST). The test set is formed by taking a $784$-point DFT of each of the $10{,}000$ test images. We obtain CV-CIFAR-10 and CV-CIFAR-100 by taking a $dN=3072$-point DFT of each of the $M=50{,}000$ training images. The test set is constructed by taking a $3072$-point DFT of each of the $10{,}000$ test images. We obtain CV-FSDD by taking a $dN=8000$-point DFT of each of the $M=2{,}700$ training audio signals. The test set is constructed by taking a $8000$-point DFT of each of the $300$ test audio signals. For CV-MNIST, CV-CIFAR-10, and CV-FSDD we have $k=10$, and for CV-CIFAR-100, we have $k=100$. RASPNet consists of $M=20{,}000$ training and $5{,}000$ test radar returns of size $dN=1680$, with $k=2$ ($2$D target localization).
We train a RVNN, CVNN, Steinmetz neural network, and analytic neural network using Cross Entropy Loss for $lN=64$ to classify the images in CV-MNIST, CV-CIFAR-10, and CV-CIFAR-100, $lN=200$ to classify the audio signals in CV-FSDD, and using MSE Loss for $lN=128$ to localize targets from the radar returns in RASPNet. These neural network architectures and hyperparameter choices are described in Section D of the Appendix, wherein we leverage the Adam optimizer (Kingma and Ba, 2014) to train each architecture using a fixed learning rate. The empirical results pertaining to this first experiment are depicted in Table 1 and Figure 3. On CV-MNIST, CV-CIFAR-10, CV-CIFAR-100, CV-FSDD, and RASPNet, we observe that the Steinmetz and analytic neural networks achieve improved generalization over the classical RVNN and the CVNN. The analytic neural network achieves the highest classification accuracy and the lowest MSE.
@@ 6.2 Channel Identification
We evaluate our proposed Steinmetz and analytic networks on the benchmark channel identification task from (Scardapane et al., 2020; Bouboulis et al., 2015). Let $\smash{X^{m}=\sqrt{1-\rho^{2}}\bar{X}\vphantom{X}^{m}+i\rho\tilde{X}\vphantom{X}^{m}}$ denote the input to the channel, wherein $\smash{\bar{X}\vphantom{X}^{m}}$ and $\smash{\tilde{X}\vphantom{X}^{m}}$ are Gaussian random variables, and $\rho$ determines the circularity of the signal. Here, $dN=5$ denotes the length of the input sequence, an embedding of the channel inputs over $dN$ time steps. The channel output, $Y^{m}\in\mathbb{C}^{k}$, is formed using a linear filter, a memoryless nonlinearity, and by adding white Gaussian noise to achieve an SNR of $5$ dB, wherein $X^{m}$ and $Y^{m}$ are equivalent to $s_{n}$ and $r_{n}$ from (Scardapane et al., 2020). We consider $M=1000$ training examples and $1000$ test examples. Each of the real-valued architectures output a $22k=$ 2D vector, $\smash{[[\hat{Y}_{R}^{m}]^{T},[\hat{Y}_{I}^{m}]^{T}]^{T}}$, where $\smash{\hat{Y}_{R}^{m},\hat{Y}_{I}^{m}\in\mathbb{R}}$, and the CVNN outputs a $k=1$-dimensional scalar, $\smash{\hat{Y}^{m}=\hat{Y}_{R}^{m}+i\hat{Y}_{I}^{m}}$, which we reshape to form $\smash{[[\hat{Y}_{R}^{m}]^{T},[\hat{Y}_{I}^{m}]^{T}]^{T}}$. We train our Steinmetz and analytic neural networks, RVNN, and CVNN using MSE Loss, minimizing the distance between $\smash{[[\hat{Y}_{R}^{m}]^{T},[\hat{Y}_{I}^{m}]^{T}]^{T}}$ and $\smash{[[Y_{R}^{m}]^{T},[Y_{I}^{m}]^{T}]^{T}}$. We select $lN=64$, and use the Adam optimizer to train each architecture using a fixed learning rate. We compute and report the MSE between the predicted and true magnitudes and phases on the test dataset in Table 2. From Table 2, we see while the magnitude prediction error is comparable between the CVNN, RVNN, and the Steinmetz and analytic neural networks, the latter pair observes a much lower phase prediction error.
@@ 7 Conclusion
In this work, we introduced Steinmetz neural networks, a new approach to processing complex-valued data using DNNs with parallel real-valued subnetworks. We provided its mathematical framework and outlined a consistency constraint to lower its generalization gap upper bound, and we presented the analytic neural network, which incorporates the consistency penalty, for practical implementation. We evaluated these networks on regression and classification benchmarks, showing improvements over existing RVNNs and CVNNs. Future work includes investigating more effective training techniques for Steinmetz neural networks and theoretical performance guarantees.
Acknowledgements
This work was supported in part by the U.S. Air Force Office of Scientific Research (AFOSR) under award FA9550-21-1-0235. Any opinions, findings and conclusions expressed in this material are those of the authors and do not necessarily reflect the views of the U.S. Department of Defense.
@@ Appendix A Consistency Constraint Derivation
As outlined in Section 4, our aim is to exploit the Steinmetz neural network architecture by deriving a consistency constraint that allows for improved control over the generalization gap, $\Delta s$. Recall Eq. 4.1, which provides an upper bound on $\Delta s$ in terms of the mutual information between $X^{m}$ and $Z^{m}$, between $Y^{m}$ and $Z^{m}$, and between $S$ and $\theta_{\psi_{S}}$. The complete derivation of this bound is provided in [Kawaguchi et al., 2023].
@@
For improved control over the generalization gap, we form a smaller upper bound on $\Delta s$ by deriving a lower bound, $\mathcal{D}(Z^{m})$, on $[I(X^{m};Z^{m})-I(Y^{m};Z^{m})+I(S;\theta_{\psi_{S}})]$. We previously stated $\mathcal{D}(Z^{m})$ is achievable by imposing a constraint on the latent representation, $Z^{m}$ (see Corollary 4.2).
@@
This result is also summarized in Theorem 4.3 of the main text. We extend this result to derive the smaller upper bound on the generalization gap, $\Delta s$, provided in Theorem 4.4 of the main text.
@@ C.2 CV-CIFAR-10 Dataset
C.3 CV-CIFAR-100 Dataset
The CIFAR-100 dataset is a collection of color images categorized into $100$ different classes, and is commonly used to train image processing systems. The CV-CIFAR-100 dataset is formed by taking a $dN=3072$-point DFT of each flattened CIFAR-100 image, yielding a $3072$-dimensional real feature vector and a $3072$-dimensional imaginary feature vector. For the classification result from Section 6.1, we consider $M=50{,}000$ training samples and $10{,}000$ test samples from CV-CIFAR-100. The features and labels within CV-CIFAR-100 are summarized as follows:
  • Target Variable: The numerical class (category) the image represents, ranging from 1 to 100 ($k=100$).
C.4 CV-FSDD Dataset
The FSDD (Free Spoken Digit Dataset) is a collection of audio recordings of spoken digits categorized into $10$ different classes, and is commonly used to train audio processing systems. The CV-FSDD dataset is formed by taking a $dN=8000$-point DFT of each flattened FSDD recording, yielding a $8000$-dimensional real feature vector and a $8000$-dimensional imaginary feature vector. For the classification result from Section 6.1, we consider $M=2{,}700$ training samples and $300$ test samples from CV-FSDD. The features and labels within CV-FSDD are summarized as follows:
  • Each feature (audio recording) is represented as a 8000-dimensional real vector and a 8000-dimensional imaginary vector, obtained by taking the DFT of the original audio signal.
  • Target Variable: The numerical digit (category) the recording represents, ranging from 1 to 10 ($k=10$).
C.5 RASPNet Dataset
The RASPNet dataset for radar adaptive signal processing consists of radar returns gathered from $100$ realistic radar scenarios located in the contiguous United States. We consider the scenario index $i=29$ from RASPNet, which corresponds to the Bonneville Salt Flats, UT, and consider the real and imaginary parts of a size $5\times 21\times 16$ radar return, comprising $5$ realizations, $21$ range bins, and $16$ channels (single pulse transmission). Accordingly, each sample is defined by a $dN=1680$-dimensional real feature vector, and a $1680$-dimensional imaginary feature vector. Each radar return contains a single point target, with position encoded in Cartesian coordinates (x, y). For the regression result from Section 6.1, we consider $M=20{,}000$ training samples and $5{,}000$ test samples from scenario $i=29$, which can be accessed here. The features and labels are summarized as follows:
@@ C.5 RASPNet Dataset
  • Target Variable: The target location in Cartesian coordinates (x, y) ($k=2$).
@@ D.3 Neural Network Training Hyperparameters
CV-CIFAR-100 No ablations Adam 0.0001 0.001
CV-FSDD No ablations Adam 0.0002 1.0e-6